• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 16:06:59 on June13,2010

    Entry number 326308

    keyword=Day Shift Summary

    Shift workers:

    Ole Hansen (SL)
    Trevor Bielarski (TO)
    Kosta Kliakhandler (3rd)

    8:00 Start shift with transverse asymmetry data taking in progress, slug #160.
    10:57 Half-wave plate OUT. Start slug #161.
    12:07 Right Q2 current goes to zero out of the blue. Cannot reset from counting house. Asking for access to reset.
    12:14 Right Q1 trips, but can be reset successfully from upstairs.
    13:00 Attempt to reset right Q2 on the power supply control panel fails. Apparently something in the supply is fried. Calling in tech on call.
    14:00 Howard (tech on call) arrives, beginning right Q2 investigation.
    15:00 ETA on magnet repair ~1hr.
    16:00 Magnet controls repair still in progress at end of shift.

    Run summary:
    Slug 160, IHWP IN
    4494 32k (beam current stepping up, might not be good)
    4495 78k
    4496 73k
    4497 88k

    Slug 161, IHWP OUT
    4498 80k
    4499 29k
    4500 12k R-Q2 died at end of run